Use a triple integral to determine the volume of the region bounded by z =


p


x


2 + y


2 and z = x


2 + y


2


In 1st octant


1
Expert's answer
2021-07-26T14:15:40-0400

"V=\int^{\pi/4}_0d\theta\int^p_0rdr\int^p_0dz=\frac{p^3\pi}{8}"